# Training Task Reasoning LLM Agents for Multi-turn Task Planning via Single-turn Reinforcement Learning

### v1

@@ 4.2 Results of Task Planning with Single-turn GRPO
Table˜1 presents the results for the four progressively challenging cooking tasks. Each table compares the proposed method against multiple baselines including different-sized Qwen models and the model without single-turn task reasoning of GRPO. These results provide strong empirical support for Theorem˜3.3, showing that single-turn GRPO training indeed leads to improved multi-turn task planning performance. Specifically, the results demonstrate the following key findings.
Significant Performance Gains by GRPO of Task Reasoning
@@
Our model with SFT and GRPO consistently outperforms much larger-size baseline models. Most notably, it achieves 0.70 success rate on the Burger task compared to 0.0 for the 14B baseline, and matches the expert’s 0.70 performance on the Cheese Burger task while being significantly more efficient (15.8 vs 17.4 average steps). The SFT-alone model provides limited success compared to the one followed by GRPO, validating the effectiveness of single-turn RL for task reasoning in Theorem˜3.3.
Efficiency and Scalability to Complex Tasks
Beyond success rates, our method consistently achieves lower average step counts for successful episodes, indicating more efficient task planning. For instance, on the Cheese Burger task, our method requires only 12.7 steps compared to 15.3 for SFT-only, demonstrating that GRPO learns more optimal policies with minimal steps of task completion in Definition˜3.2. The performance advantage is maintained even on the most challenging Double Cheese Burger task (35-step horizon), where our 1.5B model achieves 0.3 success rate while all larger LLMs fail completely.
4.3 Generalization to Unseen Task Planning
To evaluate the generalizability of LLM agent trained from single-turn task reasoning, we train LLM agents for individual tasks and evaluate their zero-shot performance on all four tasks. Tables 2, 3, and 4 present the results under each metric. These findings provide empirical validation of Corollary˜3.2 and demonstrate that our single-turn GRPO approach enables strong cross-task generalization, particularly in the direction from complex to simpler tasks, which is valuable for practical deployment scenarios where a single model needs to handle unseen varying task complexities. Some specific findings are shown below.
Consistent Generalization of Success Rate over Subtasks
Models trained on more complex tasks demonstrate better generalization to simpler tasks. The agent trained on Double Cheese Burger can solve all tasks with non-zero success rates (0.3 - 0.5), while agents trained on simpler tasks fail on more complex ones. This supports our theoretical result that complex task training captures the necessary subtask planning capabilities in Corollary˜3.2, showing that GRPO training enables generalization to subtasks with shorter horizons. We also observe that success rates generally decrease as the evaluation task becomes more complex relative to the training task, but remain positive when the evaluation task is simpler or of similar complexity.
Loss of Optimality and Failure on Complex Tasks in Generalization
Even when agents trained on complex tasks successfully generalize to simple tasks, they may require more steps than the agents trained specifically on the target simple tasks. For example, the Double Cheese Burger-trained agent requires 8.2 steps for Burger task completion compared to 8.0 steps for the Burger-trained agent, indicating some loss of efficiency and optimality during generalization. In addition, agents trained on all simpler tasks completely fail on the most complex Double Cheese Burger task, revealing the future directions of the LLM RL post-training community where complex reasoning capabilities are hard to obtain from simpler task reasoning problems.
@@ 5 Conclusion
In this paper, we present a theoretical framework demonstrating that GRPO improvements on single-turn task reasoning result in enhanced multi-turn success probability under minimal steps for challenging multi-turn interaction scenarios. Empirical results validate our approach, with our small-parameter model consistently outperforming larger baselines while requiring fewer steps for task completion. We also show cross-task generalization capabilities both theoretically and empirically. While limitations remain regarding the reliance on expert trajectories, our work shows how to leverage single-turn RL successes for complex multi-turn task planning, opening promising directions for more capable and efficient LLM-based agents.

### v2

@@ 4.2 Results of Task Planning with Single-turn GRPO
In Table˜1, the models with SFT and GRPO consistently outperform much larger-size baseline models and perform better or keep on par with expert policy. The SFT-alone model provides limited success compared to the one followed by GRPO, validating the effectiveness of single-turn RL for task reasoning in Theorem˜3.3. Beyond success rates, the model with GRPO post-training consistently achieves lower average step counts for successful episodes, demonstrating that GRPO learns more optimal policies with minimal steps of task completion in Definition˜3.2.
From Table˜2, the proposed GRPO-based approach achieves the best performance in all metrics compared to other RL post-training methods. The PPO-based agent performs the worst due to its reliance on a separate critic network for advantage estimation, which introduces additional approximation errors. Compared to critic-free RL methods like RLOO and REINFORCE++, group-relative advantage computation of GRPO provides more stable gradient estimates by normalizing within sampled responses.
4.3 Ablation Study and Generalization Analysis
Cross-task Generalization Analysis
To evaluate generalizability, we train agents on individual tasks and assess zero-shot performance across all tasks. The results in Table˜3 empirically validate Corollary˜3.2: agents trained on complex tasks generalize effectively to simpler ones, while the reverse fails. The Double Cheese Burger-trained agent achieves non-zero success rates on all tasks, whereas Burger-trained and Cheese Burger-trained agents completely fail on more complex task of Double Cheese Burger. These findings demonstrate that our single-turn GRPO approach enables meaningful cross-task generalization from complex to simpler tasks, aligning with our theoretical predictions, though highlighting that complex reasoning capabilities cannot be readily acquired from simpler task training alone.
Robustness to Noisy Non-expert Trajectories
To assess robustness to imperfect demonstrations, we evaluate the Burger task with varying ratios of noisy trajectories. Table 4 shows that our method exhibits strong robustness: even with 50% noisy data, success rate only drops to 0.50 from 0.73, while efficiency metrics (ASAT and ASST) remain stable. Performance degrades gradually as noise increases. This robustness stems from GRPO’s group-relative advantage computation, which effectively filters suboptimal demonstrations during optimization. These results indicate that our single-turn GRPO approach can successfully learn from imperfect expert trajectories, valuable for practical deployment where obtaining perfectly optimal demonstrations is challenging.
@@ 5 Conclusion
In this paper, we present a theoretical framework demonstrating that GRPO improvements on single-turn task reasoning provide a lower bound on multi-turn success probability under minimal steps for challenging multi-turn interaction scenarios. Empirical results validate our approach, with our small-parameter model consistently outperforming larger baselines while requiring fewer steps for task completion. We also show cross-task generalization capabilities both theoretically and empirically. While limitations remain regarding the reliance on expert trajectories, our work shows how to leverage single-turn RL successes for complex multi-turn task planning, opening promising directions for more capable and efficient LLM-based agents.